OpenStudy (vanillaangelic):

A tank holds 50 gallons of a mixture of acid and water that is 30% acid. If 20 gallons of acid are added to the mixture, what is the percent of acid in the new mixture? Explain step by step please~

OpenStudy (mathmate):

Hint: use \(C_1V_1 = C_2V_2+C_3V_3\) C2=0.30 V2=50 gal C3=1.0 V3=20 V1=50+20=70 Solve for C1.

OpenStudy (vanillaangelic):

I still don't understand. Can you explain it in words please?

OpenStudy (vanillaangelic):

@mathmate

OpenStudy (mathmate):

actually, the equation tells us how much pure acid there is before and after mixing. Once you have the volume of pure acid, we would divide by the volume 70 gal. to get the concentration.

OpenStudy (mathmate):

btw, CV gives the volume of the pure substance.

OpenStudy (amistre64):

parts per whole: 50(.30) + 20 is the part that is acid 50 + 20 is the whole lot of it

OpenStudy (vanillaangelic):

Thank you!! @amistre64